OpenStudy (anonymous):

It Takes 1 hour longer to fly to St.Paul at 200 mi/h than it does to return at 250 mi/h . How Far Is St.Paut ( Needs Too Be Solved In Equation Form )

OpenStudy (anonymous):

\[v=\frac{ x }{ t }\]

OpenStudy (anonymous):

d1 = d2 200 ( t +1) = 250t

OpenStudy (anonymous):

\[t=\frac{ x }{ 200 }\] \[t _{2}=\frac{ x }{ 250 }\]

OpenStudy (anonymous):

200t+200=250

OpenStudy (anonymous):

\[x/200+1=x/250\]

OpenStudy (anonymous):

200t+200=250 t @Ambbiiee Can you solve for t ?

OpenStudy (anonymous):

my logic is not corrct is it @Chlorophyll

OpenStudy (anonymous):

200=250 -200 -200 _____________ 200t =50

OpenStudy (anonymous):

I'm not sure, just see if we come to the same result!

OpenStudy (anonymous):

t=4

OpenStudy (anonymous):

Now plug back to d1 = 200t

OpenStudy (anonymous):

i get it now thankyou !(:

OpenStudy (anonymous):

I mean plug it back to 250t => d2 = 1,000 mi